# BoxDreamer: Dreaming Box Corners for Generalizable Object Pose Estimation

### v1

@@ Abstract
This paper presents a generalizable RGB-based approach for object pose estimation, specifically designed to address challenges in sparse-view settings. While existing methods can estimate the poses of unseen objects, their generalization ability remains limited in scenarios involving occlusions and sparse reference views, restricting their real-world applicability. To overcome these limitations, we introduce corner points of the object bounding box as an intermediate representation of the object pose. The 3D object corners can be reliably recovered from sparse input views, while the 2D corner points in the target view are estimated through a novel reference-based point synthesizer, which works well even in scenarios involving occlusions. As object semantic points, object corners naturally establish 2D-3D correspondences for object pose estimation with a PnP algorithm. Extensive experiments on the YCB-Video and Occluded-LINEMOD datasets show that our approach outperforms state-of-the-art methods, highlighting the effectiveness of the proposed representation and significantly enhancing the generalization capabilities of object pose estimation, which is crucial for real-world applications. The code will be released at project page.
@@ 1 Introduction
However, for sparse-view scenarios, retrieving reference views with similar viewpoints is challenging, leading to poor pose initialization. The poor pose initialization adversely affects subsequent modules that refine the pose. Moreover, for objects in query images with severe occlusion, selecting the correct reference is highly challenging, often leading to the failure of the pose estimation process.
@@
Concretely, our approach involves two main steps: recovering the 3D bounding box and predicting 2D projections of the 3D bounding box for the query image. First, we use sparse-view reconstruction tools to estimate camera poses and recover the object’s approximate structure to compute its 3D bounding box. To leverage 3D corner locations for object pose estimation, we project the 3D bounding box corners onto each reference image to generate dense projection maps, which are then fed into an end-to-end transformer decoder to predict the 2D projection of the 3D bounding box corners in the query view. The global nature of object box corners enables our method to infer box corner positions based on the object’s visible parts and reference demonstration. Finally, the 6DoF object pose is recovered using a Perspective-n-Point (PnP) algorithm with the 2D-3D correspondences. Note that our method does not require an accurate 3D bounding box for object pose estimation, as demonstrated by our experimental results in Sec. 4.5.
@@
  • The first to propose object bounding box corners as an intermediate representation for practical object pose estimation.
@@ 2.1 CAD-Model-Based Object Pose Estimation
Model-based object pose estimation methods can be broadly classified into two categories. Instance-level pose estimators [34, 44, 32, 22, 35, 48, 5, 28, 23, 18, 15] require separate training for each object and often rely on CAD models either to establish 2D–3D correspondences [32, 34, 35, 5, 28, 15] or for rendering and pose refinement [22, 48]. There are also methods that attempt to regress object pose from RGB images directly. [44, 23, 18]. Although they achieve high accuracy for known objects, they are not generalizable to unseen objects.
@@ 2.3 Object 3D Representation
Providing 3D information about the object is essential for generalizable object pose estimation. The OnePose family employs dense point clouds to represent objects, while Gen6D constructs a volumetric representation from multi-view images and utilizes it for pose refinement. LatentFusion [33] establishes 3D latent to represent objects for pose estimation and refinement. GS-Pose [2] represents objects using 3D Gaussians [19] and performs pose refinement based on it. Despite their advancements, none of these methods offer a simple yet robust 3D representation for 6-DoF pose estimation for any unseen object, as they heavily rely on dense reference images or struggle under challenging conditions such as occlusion or textureless surfaces. To the best of our knowledge, we are the first to propose using the bounding box corners as a compact and effective representation for generalizable object pose estimation.
@@ 2.4 Sparse-View Object Pose Estimation
Some recent works have addressed sparse-view pose estimation [26, 27, 38]. For example,  [26] employs an Object Gaussian representation with a render-and-refine strategy, achieving high accuracy but requiring about 10 minutes to construct the representation, which limits practical use. Meanwhile,  [27] generates novel view embeddings, and  [38] leverages a novel view synthesis (NVS) model; however, both methods rely on generative performance, which may face challenges on real-world diverse objects, and NVS is also time-consuming. FoundationPose [43] shows impressive performance with sparse views through large-scale training on realistic synthetic datasets, yet its dependence on depth input restricts real-world use. In contrast, our method uses lightweight and effective representations for robust, real-time pose estimation from sparse-view references.
@@ 3 Method
To this end, we utilize an any off-the-shelf reconstruction method to obtain the point cloud $\mathbf{P}$ of the target object, along with its 3D bounding box $\mathbf{B}$ and corresponding reference poses $\{\mathbf{\xi}_{0},\dots,\mathbf{\xi}_{i}\}$ (see Sec. 3.1).
Then, we propose a novel bounding box estimation network to predict the projection $\mathbf{{b}}_{q}$ of the query view’s object bounding box. Based on these predictions, we establish 2D-3D correspondences between the predicted 2D box corners $\mathbf{b}_{q}$ and the reconstructed 3D box corners $\mathbf{B}$, and recover the query object pose $\mathbf{\xi}_{q}$ using the PnP algorithm (see Sec. 3.2).
@@ 3.1 Preparation of Object Bounding Box
3D Bounding Box as object representation. To address the limitations of OnePose, we remove the reliance on dense point clouds. Instead, we represent the object with a 3D bounding box $\mathbf{B}$, which serves as a rough spatial prior for pose estimation. The advantage of using a bounding box rather than a dense point cloud is that 3D bounding boxes can be restored from sparse view inputs, facilitating object coordinate recovery even when dense observations are unavailable.
To obtain the 3D bounding box $\mathbf{B}$, we first collect a sparse set of reference images. Using these images, we apply an off-the-shelf reconstruction method [42, 21, 40, 46] to extract the object’s sparse geometry. We then filter out irrelevant points according to the object detection results, ultimately obtaining the 3D bounding box $\mathbf{B}$ that encloses the object.
By providing cropped reference images to fully feed-forward reconstruction methods, we obtain both the object point cloud $\mathbf{P}$ and the camera poses $\{\mathbf{\xi}_{0},\dots,\mathbf{\xi}_{i}\}$ for the reference views. Specifically, for methods like DUSt3R [42] (which we used in experiments), sparse-view images are directly fed into the network to predict a Pointmap, which establishes dense correspondences between image pixels and associated 3D points. Once the Pointmap is obtained, the reference camera poses are efficiently recovered by applying the Perspective-n-Point (PnP) algorithm, leveraging the 2D-3D correspondences provided by the Pointmap. This process simultaneously recovers both 3D geometry and camera poses, making it particularly effective in sparse-view scenarios.
@@
2D Heatmap as box representation.
We observed that directly using the eight corner coordinates of a 3D bounding box often degrades performance due to the inherent sparsity of the input signals. To better leverage the reference information, we project the 3D bounding box onto the image plane and generate a heatmap representing its 2D spatial distribution. This strategy aligns well with vision transformers, which can effectively process heatmap data. Specifically, given the eight 3D corners of the bounding box $\mathbf{B}$, we compute their projection onto the image plane $\mathbf{b}$ via the perspective projection function.
Directly using a one-hot heatmap for these eight corners tends to yield supervision signals that are excessively sparse and insufficiently smooth, which can hamper model learning. To mitigate this, we draw inspiration from CornerNet [20], which applies Gaussian smoothing around each ground-truth corner location. This approach eases the penalty for negative locations within a certain radius and avoids sharp transitions between positive and negative regions.
In practice, we find CornerNet’s hyperparameters not directly transferable to our setting. Thus, we redefine the heatmap function as:
@@
Where the squared distance is replaced by the distance itself, further enhancing heatmap smoothness. We set the radius $\sigma$ to one-tenth of the object size, defined as the average squared pixel distance from the eight corners to their center.
@@ 3.2 Object Pose Estimation
For the query image, rather than heatmap tokens, we use learned query tokens $\mathbf{Q}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times d}$. After these preparations, we flatten the reference and query features and concatenate them along the patch dimension, producing a 1D token sequence with length $l=(N+1)\times\frac{H\times W}{p^{2}}$, where $N$ is the number of reference views. This sequence is then fed into a full self-attention transformer decoder comprising $L$ layers, generating the query bounding box features $\mathbf{F}_{q}^{\prime}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times d}$.
Finally, we apply a linear layer to map the query bounding box features back to the original heatmap dimension and unpatchify the result to obtain the final query bounding box heatmap:
@@
Given the predicted query bounding box heatmap and the reconstructed 3D bounding box, we can establish 2D-3D correspondences, subsequently recovering the object pose $\mathbf{\xi}_{q}$ using the Perspective-n-Point (PnP) algorithm.
@@ 3.3 Training
Supervision. We use Smooth L1 Loss [11] to supervise both the predicted bounding box heatmap (coarse) and each specific corner point (fine). Through our smoother heatmap settings.
@@ 3.4 Implementation Details
Data Augmentation. To boost the generalizability, we randomly rotate the 3D bounding box (within $[-\pi,\pi]$ along a random axis) to break its direct association with semantic information. We also apply RGB augmentations such as motion blur and noise, and composite synthetic images with random SUN2012 [45] backgrounds. Moreover, we further improve occlusion handling by randomly occluding objects through truncation, masking, or inserting images of other objects.
Network Architecture. The network leverages a transformer decoder consisting of $L=12$ layers, $d=768$ hidden units, and $8$ attention heads. The patch size is set to $14$ following the DINOv2-base [29] model.
Training. The model is trained using the AdamW optimizer with an initial learning rate of $10^{-4}$, following a cosine decay schedule. During training, the number of reference images is dynamically sampled between $1$ and $15$, while the batch size per GPU varies from $144$ to $18$. The training process spans 100 epochs and is conducted on 8 A100-SXM4-80GB GPUs.
@@ 4.1 Experimental Setup and Baselines
  • LINEMOD and Occluded LINEMOD: Adopt the standard training-testing split from prior studies [34, 25, 36, 12], providing approximately 180 reference images per object.
  • YCB-Video: We create three distinct reference databases to assess robustness across varying reference qualities: (a) Sparse Database: Consists of 16 reference images per object sampled from different video sequences, as proposed in FoundationPose [43]. (b) Most-overlapping Database: To best test baseline methods, we build this database by selecting the most overlapping video sequence as the reference database. (c) Occlusion-minimizing Database: To avoid occlusion issues in the most overlapping sequence, which introduces low-quality reference views, we manually choose a reference sequence that maintains high overlap while ensuring minimal occlusion. Details on the construction of the YCB-Video reference databases are provided in the supplementary material.
  • OnePose and OnePose-Lowtexture: Use the first sequence as the reference database following the settings in [12] and  [36].
To ensure consistency across all datasets and methods, we uniformly sample sparse-view references using the Farthest Point Sampling (FPS) algorithm. For databases with more than five reference images, unless stated otherwise, BoxDreamer initially selects five neighbors by directly computing the cosine similarity between the query and the reference DINOv2 features.
@@
4.2 Results on LINEMOD and Occluded LINEMOD
@@
Table 1: Comparison on the LINEMOD subset. ADD(s)-0.1d is reported. Gen6D indicates that all reference images are used for refinement, and Gen6D is the public results from GS-Pose, GS-Poseinit reports the initial results without Gaussian Splatting refinement.
@@
Table 2: Comparison on the Occluded LINEMOD dataset. ADD(s)-0.1d and Proj-2d@5px are reported, indicates provide ground-truth detection results for Gen6D and objects in italic are included in the Gen6D training set.
@@ 4.3 Results on YCB-Video
Table 3: Performance of the YCB-Video dataset on Sparse Database. indicates provide ground-truth detection results for Gen6D. indicates the background has been removed based on the ground-truth mask to help Gen6D to achieve better detection results.
@@
Table 4: Average Performance on YCB-Video dataset with different databases. ADD and ADD-S are reported. Gen6D indicates that the method benefits from ground-truth detection results.
4.4 Results OnePose-Lowtexture
@@
Table 5: Performance comparison on OnePose-LowTexture dataset. ADD-0.1d is reported.
@@ 4.5 Analysis
Figure 4: Effects of the number of reference images.
Affects of Noisy 3D Bounding Boxes. We use DUSt3R to recover the object bounding boxes by filtering points with multiview detections. This process can introduce noise in the recovered boxes, so we evaluate our method’s robustness to variations in bounding box quality. As shown in Fig. 5, our approach remains robust across different bounding boxes. Furthermore, experiments on the LINEMOD dataset with various bounding box sources (see Table 6) confirm the method’s adaptation to noisy box estimates.
@@
Table 6: Performance on LINEMOD using different bounding box sources.
@@
Table 7: Effects of Real Data Training.
Running time. On an Intel i7-13700KF CPU and NVIDIA RTX 4090 GPU, our method processes a query image in approximately 23 ms using 10 reference images for pose inference. Specifically, it takes about 6 ms for the DINOv2 encoder, 12 ms for the bounding box decoder, and 0.9 ms for pose recovery, confirming its real-time performance. Prior to inference, reconstructing 10 reference images with DUSt3R (including model loading, reconstruction, and point cloud filtering) requires roughly 11 s.

### v2

@@ Abstract
This paper presents a generalizable RGB-based approach for object pose estimation, specifically designed to address challenges in sparse-view settings. While existing methods can estimate the poses of unseen objects, their generalization ability remains limited in scenarios involving occlusions and sparse reference views, restricting their real-world applicability. To overcome these limitations, we introduce corner points of the object bounding box as an intermediate representation of the object pose. The 3D object corners can be reliably recovered from sparse input views, while the 2D corner points in the target view are estimated through a novel reference-based point synthesizer, which works well even in scenarios involving occlusions. As object semantic points, object corners naturally establish 2D-3D correspondences for object pose estimation with a PnP algorithm. Extensive experiments on the YCB-Video and Occluded-LINEMOD datasets show that our approach outperforms state-of-the-art methods, highlighting the effectiveness of the proposed representation and significantly enhancing the generalization capabilities of object pose estimation, which is crucial for real-world applications.
@@ 1 Introduction
However, for sparse-view scenarios, retrieving reference views with similar viewpoints is challenging, leading to inaccurate pose initialization. The suboptimal pose initialization impairs subsequent modules that refine the pose. Moreover, for objects in query images with severe occlusion, selecting the correct reference is highly challenging, often leading to the failure of the pose estimation process.
@@
Concretely, our approach involves two main steps: recovering the 3D bounding box and predicting 2D projections of the 3D bounding box corners for the query image. First, we use sparse-view reconstruction tools to estimate camera poses and recover the object’s approximate structure to compute its 3D bounding box. To leverage 3D corner locations for object pose estimation, we project the 3D bounding box corners onto each reference image to generate corner heatmap, which are then fed into an end-to-end transformer decoder to predict the 2D projection of the 3D bounding box corners in the query view. The global nature of object box corners enables our method to infer box corner positions based on the object’s visible parts and reference demonstration. Finally, the 6DoF object pose is recovered using a Perspective-n-Point (PnP) algorithm with the 2D-3D correspondences. Note that our method does not require an accurate 3D bounding box for object pose estimation, as demonstrated by our experimental results in Sec. 4.5.
@@
  • We propose object bounding box corners as an intermediate representation for generalizable object pose estimation.
@@ 2.1 CAD-Model-Based Object Pose Estimation
Model-based object pose estimation methods can be broadly classified into two categories. Instance-level pose estimators [33, 43, 31, 22, 34, 47, 5, 28, 23, 18, 15] require separate training for each object and often rely on CAD models either to establish 2D–3xD correspondences [31, 33, 34, 5, 28, 15] or for rendering and pose refinement [22, 47]. There are also methods that attempt to regress object pose from RGB images directly. [43, 23, 18]. Although they achieve high accuracy for known objects, they are not generalizable to unseen objects.
@@ 2.3 Object 3D Representation
Providing 3D information about the object is essential for generalizable object pose estimation. The OnePose family employs dense point clouds to represent objects, while Gen6D constructs a volumetric representation from multi-view images and utilizes it for pose refinement. LatentFusion [32] establishes 3D latent to represent objects for pose estimation and refinement. GS-Pose [2] represents objects using 3D Gaussians [19] and performs pose refinement based on it. Despite their advancements, none of these methods offer a simple yet robust 3D representation for 6-DoF pose estimation for any unseen object, as they heavily rely on dense reference images or struggle under challenging conditions such as occlusion or textureless surfaces. In contrast, our approach innovatively uses bounding box corners as a compact and effective representation for generalizable object pose estimation.
@@ 2.4 Sparse-View Object Pose Estimation
Some recent works have addressed sparse-view pose estimation [26, 27, 37]. For example, Luo et al. [26] employs an Object Gaussian representation with a render-and-refine strategy, achieving high accuracy but requiring about 10 minutes to construct the representation, which limits practical use. Meanwhile, Nguyen et al. [27] generates novel view embeddings, and Sun et al. [37] leverages a novel view synthesis (NVS) model; however, both methods rely on generative performance, which may face challenges on real-world diverse objects, and NVS is also time-consuming. FoundationPose [42] shows impressive performance with sparse views through large-scale training on realistic synthetic datasets, yet its dependence on depth input restricts real-world use. In contrast, our method uses lightweight and effective representations for robust, real-time pose estimation from sparse-view references.
@@ 3 Method
To this end, we utilize an off-the-shelf reconstruction method to obtain the point cloud $\mathbf{P}$ of the target object, along with its 3D bounding box $\mathbf{B}$ and corresponding reference poses $\{\mathbf{\xi}_{0},\dots,\mathbf{\xi}_{i}\}$ (see Sec. 3.1).
Then, we propose a novel bounding box estimation network to predict the projected box corners $\mathbf{b}_{q}$ in the query view. Based on these predictions, we first establish 2D-3D correspondences between the predicted 2D corners $\mathbf{b}_{q}$ and the reconstructed 3D corners $\mathbf{B}$, then recover the query object pose $\mathbf{\xi}_{q}$ using the PnP algorithm (see Sec. 3.2).
@@ 3.1 Preparation of Object Bounding Box
3D Bounding Box as object representation. To address the limitations of OnePose, we avoid reliance on dense point clouds. Instead, we represent the object with a 3D bounding box $\mathbf{B}$, which serves as a rough spatial prior for pose estimation. The advantage of using a bounding box rather than a dense point cloud is that 3D bounding boxes can be restored from sparse view inputs, facilitating object coordinate recovery when dense observations are unavailable.
To obtain the 3D bounding box $\mathbf{B}$, we first collect sparse reference images. Using these images, we apply an off-the-shelf reconstruction method [41, 21, 39, 45] to extract the object’s sparse geometry. We then filter out irrelevant points according to the object detection results, ultimately obtaining the 3D bounding box $\mathbf{B}$ that encloses the object.
By feeding cropped reference images to fully feed-forward reconstruction methods, we obtain the object point cloud $\mathbf{P}$ and the camera poses $\{\mathbf{\xi}_{0},\dots,\mathbf{\xi}_{i}\}$ for the reference views. Specifically, for methods like DUSt3R [41] (used in our experiments), sparse-view images are directly fed into the network to predict Pointmaps, which establishes dense correspondences between image pixels and associated 3D points. Once the Pointmap is obtained, the reference camera poses are efficiently recovered by applying the Perspective-n-Point (PnP) algorithm, leveraging the 2D-3D correspondences provided by the Pointmap. This process simultaneously recovers both 3D geometry and camera poses, making it particularly effective in sparse-view scenarios.
@@
2D Heatmap as box representation. We observed that the direct use of the eight corner coordinates of a 3D bounding box often degrades performance due to the inherent sparsity of the input signals. To better leverage the reference information, we project the 3D bounding box onto the image plane and generate a heatmap representing its 2D spatial distribution. This strategy aligns well with vision transformers, which can effectively process heatmap data. Specifically, given the eight 3D corners of the bounding box $\mathbf{B}$, we compute their projection onto the image plane $\mathbf{b}$ via the perspective projection function.
Nevertheless, using a one-hot heatmap for these projected corners yields supervision signals that are excessively sparse and insufficiently smooth, which can hamper model learning. To mitigate this, we draw inspiration from CornerNet [20], which applies Gaussian smoothing around each ground-truth corner. This approach eases the penalty for negative locations within a certain radius and avoids sharp transitions between positive and negative regions.
In practice, we observed that the hyperparameters from CornerNet were not directly transferable to our setting. To further enhance the smoothness of the heatmap, we redefine the heatmap function as:
@@
We set the denominator term $2\sigma^{2}$ to the square of one-tenth of the object size, which we define for each corner $i$ as its pixel distance to the object’s 2D center.
@@ 3.2 Object Pose Estimation
For the query image, rather than heatmap tokens, we use learned query tokens $\mathbf{Q}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times d}$. We then flatten the reference and query features and concatenate them along the patch dimension, producing a 1D token sequence with length $l=(N+1)\times\frac{H\times W}{p^{2}}$, where $N$ is the number of reference views. This sequence is then fed into a full self-attention transformer decoder comprising $L$ layers, generating the query bounding box features $\mathbf{F}_{q}^{\prime}\in\mathbb{R}^{\frac{H}{p}\times\frac{W}{p}\times d}$.
Finally, we apply a linear layer to map the query features back to the original heatmap dimension and unpatchify the result to obtain the final query bounding box heatmap:
@@
Given the predicted query bounding box heatmap and the reconstructed 3D bounding box, we establish 2D-3D correspondences by assigning each heatmap channel to a specific bounding box corner according to a predefined channel-wise order. We then recover the object pose $\mathbf{\xi}_{q}$ using the Perspective-n-Point (PnP) algorithm.
@@ 3.3 Training
Supervision. We use Smooth L1 Loss [11] to supervise both the predicted bounding box heatmap (coarse) and each specific corner point (fine).
@@ 3.4 Implementation Details
Data Augmentation. To boost the generalizability, we randomly rotate the 3D bounding box (within $\pi,\pi]$ along a random axis) to break its direct association with semantic information. We also apply RGB augmentations such as motion blur and noise, and composite synthetic images with random SUN2012 [44] backgrounds. Moreover, we further improve occlusion handling by randomly occluding objects through truncation and masking.
Network Architecture. The network leverages a transformer decoder consisting of $L=12$ layers, $d=768$ hidden size, and $8$ attention heads. The patch size is set to $14$ following the DINOv2-base [29] model.
Training. The model is trained using the AdamW optimizer with an initial learning rate of $10^{-4}$, following a cosine decay schedule. During training, the number of reference images is dynamically sampled between $1$ and $15$, while the batch size per GPU varies from $144$ to $18$. The model is trained for 100 epochs using 8 A100-SXM4-80GB GPUs.
@@ 4.1 Experimental Setup and Baselines
  • LINEMOD and Occluded LINEMOD: We use the standard training-testing split from prior studies [33, 25, 35, 12], with $\sim$180 reference images per object.
  • YCB-Video: We create three distinct reference databases to assess robustness across varying reference qualities: (a) Sparse Database: Contains 16 reference images per object, sampled from distinct video sequences as in FoundationPose [42]. (b) Most-overlapping Database: To best test baseline methods, we build this database by selecting the most overlapping video sequence as the reference database. (c) Occlusion-minimizing Database: To mitigate occlusion in the most overlapping sequence, we manually select a reference sequence that maintains high overlap with minimal occlusion. Details on the construction of the YCB-Video reference databases are provided in the supplementary material.
  • OnePose and OnePose-Lowtexture: We use the first sequence as the reference database, as in [35, 12].
To ensure consistency across all datasets and methods, we uniformly sample sparse-view references using the Farthest Point Sampling (FPS) algorithm. For experiments with more than five reference images, unless stated otherwise, BoxDreamer initially selects five neighbors by directly computing the cosine similarity between the query and the reference DINOv2 features.
@@
4.2 Results on LM and Occluded LM
@@
Table 1: Comparison on the LINEMOD subset. ADD(s)-0.1d is reported. Gen6D indicates that all reference images are used for refinement, and Gen6D is the public results from GS-Pose, GS-Poseinit reports the initial results without Gaussian Splatting refinement. We’ve highlighted the best, second-best, and third-best performing metrics in the table.
@@
Table 2: Comparison on the Occluded LINEMOD dataset. ADD(s)-0.1d and Proj-2d@5px are reported. Indicates that ground-truth detection results were provided for Gen6D. Objects in italic were included in the Gen6D training set. We’ve highlighted the best and second-best and third-best results.
@@ 4.3 Results on YCB-Video
Table 3: Performance of the YCB-Video dataset on Sparse Database. Indicates that ground-truth detection results were provided for Gen6D. Indicates that the background was removed using ground-truth masks to improve Gen6D’s detection results. We’ve highlighted the best, second-best, and third-best results.
@@
Table 4: Average Performance on YCB-Video dataset with different reference databases. ADD and ADD-S are reported. Gen6D indicates that the method benefits from ground-truth detection results. The best results are shown in bold.
4.4 Results on OnePose-LowTexture
@@
Table 5: Performance comparison on OnePose-LowTexture dataset. ADD-0.1d is reported. We’ve highlighted the best, second-best, and third-best results.
@@ 4.5 Analysis
Figure 4: The trend of performance on the LINEMOD dataset as the number of reference views changes.
Effects of Noisy 3D Bounding Boxes. We use DUSt3R to recover the object bounding boxes by filtering points with multiview detections. This process can introduce noise in the recovered boxes, so we evaluate our method’s robustness to variations in bounding box quality. As shown in Fig. 5, our approach remains robust across different bounding boxes. Furthermore, experiments on the LINEMOD dataset with various bounding box sources (see Table 6) confirm the method’s adaptation to noisy box estimates.
@@
Table 6: Performance on LINEMOD using different bounding box sources. The best results are shown in bold.
@@
Table 7: Performance on LINEMOD with different training datasets. The best results are shown in bold.
Running time. On an Intel i7-13700KF CPU and NVIDIA RTX 4090 GPU, our method processes a query image in approximately 17 ms using 5 reference images for pose inference. Specifically, it takes about 5.7 ms for the DINOv2 encoder, 11.1 ms for the bounding box decoder, and 0.15 ms for pose recovery, confirming its real-time performance. Prior to inference, reconstructing 5 reference images with DUSt3R (including model loading, reconstruction, and point cloud filtering) requires roughly 5.3 s. This process is conducted offline for once and does not affect real-time inference.